OpenStudy (adi3):

Will medal Please help. consider f(x) = 2x+1/x+3. Find f^-1(x)

OpenStudy (adi3):

@dan815

OpenStudy (jhannybean):

So first you want to set your function as y =...

OpenStudy (adi3):

ok so 2y+1/y+3 = x

OpenStudy (adi3):

then

OpenStudy (jhannybean):

Alright now multiply x by y+3 to create a linear function.

OpenStudy (jhannybean):

What do you get with that.

OpenStudy (adi3):

xy+3x

OpenStudy (jhannybean):

\[x=\frac{2y+1}{y+3} \qquad xy +3x = 2y+1\] Do you see how this is working out?

OpenStudy (adi3):

xy+3x=2y+1, how did you do this

OpenStudy (adi3):

ohh did u cross muliply

OpenStudy (jhannybean):

\[(y+3) \cdot x= \frac{2y+1}{y+3}\cdot (y+3) \]

OpenStudy (jhannybean):

Yes.

OpenStudy (adi3):

ok, then

OpenStudy (jhannybean):

Then subtract -2y from both sides of the equation. \[xy-2y+3x=1\] And then youd subtract -3x from both sides of the equation. \[xy-2y=1-3x\]

OpenStudy (jhannybean):

Following?

OpenStudy (adi3):

why did -3x switch?

OpenStudy (jhannybean):

because we want to isolate all the terms with `y` to the left side of the equation, and 3x does not have a `y` so we send it to the other side. In the process of solving for inverse functions, we switch x and y, then resolve for y.

OpenStudy (adi3):

ohhh ok, what to do next

OpenStudy (jhannybean):

Now we factor out a y from the left hand side of the equation. What do you end up with when you do that?

OpenStudy (adi3):

y(x-2)

OpenStudy (jhannybean):

Good.

OpenStudy (jhannybean):

now we divide both sides of the equation by (x-2) to solve for y. what is your equation then?

OpenStudy (adi3):

1-3x/x+2

OpenStudy (jhannybean):

careful with your signs. When you divide both sides by (x-2) it doesnt change to (x+2)

OpenStudy (jhannybean):

\[y(x-2) = 1-3x\qquad \implies \qquad y=\frac{1-3x}{x-2}\]

OpenStudy (adi3):

Thank you.

OpenStudy (jhannybean):

now all you have to do is rewrite y as f\(^{-1}\)(x) and this is your inverse function.